# Quantum Energy Teleportation versus Information Teleportation

### v1

@@ Abstract
Quantum energy teleportation (QET) is the phenomenon in which locally inaccessible energy is activated as extractable work through collaborative local operations and classical communication (LOCC) with an entangled partner. It closely resembles the more well-known quantum information teleportation (QIT) where quantum information can be sent through an entangled pair with LOCC. It is tempting to ask how QET is related to QIT. Here we report a first study of this connection. Despite the apparent similarity, we show that these two phenomena are not only distinct but moreover are mutually exclusive to each other. We show a perturbative trade-off relation between their performance in a thermal entangled chaotic many-body system, in which both QET and QIT are simultaneously implemented through a traversable wormhole in an emergent spacetime. To better understand their competition, we study the finite-dimensional counterpart of two entangled qudits and prove a universal non-perturbative trade-off bound. It shows that for any teleportation scheme, the overall performance of QET and QIT together is constrained by the amount of the entanglement resource. We discuss some explanations of our results.
@@
To answer this question, we examine a setup where QIT and QET are simultaneously implemented with a single protocol. It concerns two entangled strongly coupled many-body quantum systems whose complicated dynamics can be more easily understood with the help of an emergent spacetime [10, 11, 12]. For a thermal field double state entangling two such systems, a special family of information teleportation protocols can be described through the physics of a traversable wormhole in the emergent spacetime [13, 14]. This has inspired a series of works exploring the deep connection between teleportation and wormholes [15, 16, 17, 18, 19, 20, 21, 22, 23].
We observe that the same protocol also simultaneously teleports energy. Furthermore, to our surprise, we find that QET and QIT are competitive with each other in this example. Their exclusivity is manifest in a perturbative trade-off relation that we demonstrate.
Motivated by this observation, we look for universal non-perturbative trade-off relations between QET and QIT. To this end, we resort to the finite-dimensional counterpart of the traversable wormhole setup that consists of two arbitrarily entangled systems with some arbitrary decoupled Hamiltonian. We prove a universal non-perturbative trade-off bound for both the thermal-entangled states and arbitrary entangled states, showing that the overall performance of QET and QIT altogether is upper-bounded by the amount of shared entanglement. The bounds are universal and non-perturbative in the sense that they apply to any measurement schemes of Alice and any conditional unitaries of Bob.
Trade-off in traversable wormholes.—We find that a setup that naturally incorporates both teleportations is the traversable wormhole in holographic systems [13, 14]. It was discovered to be a convenient gravitational description of QIT. We use a concrete example of such systems in AdS2/CFT1 correspondence [14] to explain the result. In the bulk, it is described by Jackiw–Teitelboim (JT) gravity, and it is dual to a Schwarzian quantum mechanics theory on the boundary. Technically, Schwarzian quantum mechanics has a continuous energy spectrum. However, the Schwarzian describes the low energy dynamics of the Sachdev-Ye-Kitaev (SYK) model at the large $N$ limit [10], so we are allowed to think of the boundary dual of JT gravity as a normal quantum mechanical system like the SYK model with $N$ fermions.
We consider a two-sided black hole with inverse temperature $\beta$, which is dual to the thermal field double state [24],
@@
The wormhole in Fig. 1 is a priori not traversable, as one can read from the causal structure depicted by the diagonal dotted lines in the Penrose diagram. Any two points between the left and right Rindler wedges are space-like separated. On the boundary, this indicates an entangled state alone can not be used for communication. However, seminal works [13, 14] point out that the wormhole can be made traversable by introducing a double trace deformation on two sides $e^{igV_{AB}}$, where $V_{AB}:=\frac{1}{K}\sum_{i=1}^{K}O_{i}^{A}O_{i}^{B}$. This interaction couples $K$ pairs of simple operators $O_{i}$ on $\mathcal{O}(1)$ of fermions and $g$ controls the strength. The deformation with $g>0$ creates an ingoing shock wave with negative null energy that moves the black hole horizons inward, rendering a traversable wormhole. Namely, a particle carrying information thrown in the past from Alice’s side can now make its way to Bob’s side. For this description to be accurate, one technically chooses $K$ to be large $K\sim\mathcal{O}(N)$ so as to have a sharp bulk dual as a classical shock-wave geometry.
As far as Bob’s system is concerned, one can view the effect of the double trace deformation $e^{igV_{AB}}$ on Bob’s system, as an LOCC protocol $\prod_{x}\Pi_{x}^{A}\otimes U^{B}_{x}$, in which Alice measures her share of the entanglement pair and Bob implements a conditional unitary on his share upon receiving the measurement result from Alice after a time delay of $t$ (cf. Fig. 1 and Supplemental Materials for details). Usually, $e^{igV_{AB}}$ is acted on the time-symmetric slice so $t=0$, but we consider acting it asymmetrically to fit the operational interpretation as an LOCC, so a delay of $t>0$ is inevitable. In this protocol, Alice can transmit information by inserting a simple operator with a species index at around the scrambling time $T$ before the measurement. The information will be transmitted to Bob’s side after Bob’s unitary decoding. After a scrambling time, the information will re-focus onto simple operators. (cf. [17] for details in the context of the SYK model.)
The first observation we make is that the traversable wormhole protocol actually also implements quantum energy teleportation. Without Alice’s communication, Bob cannot extract energy by doing a unitary, because Bob’s reduced density matrix is a thermal density matrix. With Alice’s measurement result $x$, Bob applies the conditional unitary decoder $U_{x}$ after some communication delay $t$. By energy conservation, Bob then gains energy that amounts to the negative energy that is injected into his system via the conditional unitary. In the Supplemental Materials, we calculate this teleported energy to be
@@
We can make this trade-off more precise by quantifying the teleported information. We use $k$ different particle species indexed by $x\in[k]$ to carry the message. To calculate how much information can be teleported, we consider sending a maximally mixed message. By linearity, the state that Bob receives is $\rho=\frac{1}{k}\sum_{x=1}^{k}\rho_{x}$, where $\rho_{x}$ is the protocol output for the message $x$. The entropy of the message is measured by $\Delta S:=S(B)_{\rho}-S(B)_{\rho_{0}}$, where $\rho_{0}$ is Bob’s density matrix after the double trace deformation but without the message particle insertion  [^1] One could use other measures such as the coherent (Holevo) information to quantify the quantum (classical) information capacity. They all behave very similarly to the subtracted entropy w.r.t. varying k [26, 27]. We therefore use the simplest figure of merit as a proof of principle..
We can calculate each term in $\Delta S$ using the (quantum) Ryu-Takayanagi formula [28, 29]. The state $\rho_{0}$ describes a bulk without matter and its entropy is given by $S(B)_{\rho_{0}}=A/4G_{N}$, where $A$ is the area of the shifted bifurcation surface after the deformation. On the other hand, $S(B)_{\rho}=A/4G_{N}+S_{\mathrm{matter}}$  [^2] Here the area term will be the same as for $\rho_{0}$ under the approximation of the messenger particle being light., where $S_{\mathrm{matter}}$ is the quantum field theoretic entropy of the message particle. We know that $S_{\mathrm{matter}}$ of a uniform mixture of $k$ particle species grows like $\log k$ until it saturates at the Bekenstein bound set by $\beta\delta E$, and $\delta E:=\langle H_{B}\rangle_{\rho}-\langle H_{B}\rangle_{\rho_{0}}$. Hence, $\Delta S$ increases with $\log k$, but eventually caps off due to the Bekenstein bound [25, 30, 31, 32]. The maximal $\Delta S(B)$ thus measures the capacity of QIT  [^3] This calculation is explicitly done using the replica trick in the upcoming paper by Lu-Yang-Zheng [33]..
Therefore, the loss in QET is totally compensated by the gain in QIT  [^4] As always in thermodynamics, the (inverse) temperature $\beta$ here is the conversion factor that is necessary for comparing the distinct resources of information and energy.. In other words, upon varying the relevant parameters of the protocol, such as the conformal weight (mass) of the message particle $m$, the total variation of the combined QET+QIT figure of merit $\Delta S+\beta E$ is zero.
$\delta(\Delta S+\beta E)=\delta\Delta S+\beta\delta E=0\ .$ (3)
@@
There are some drawbacks of the trade-off (3). It is only perturbative because we cannot vary $m$ too much, as it causes significant backreaction rendering our semiclassical analysis inaccurate. Moreover, it only applies to a specific family of LOCC schemes that admit the bulk description as traversable wormholes. In seeking of a non-perturbative trade-off relation that applies to all LOCC protocols, we resort to a finite-dimensional counterpart.
Trade-off in finite dimensions.—Consider a finite-dimensional thermal field double state (1), $|\tau_{\beta}\rangle_{AB}$, between Alice and Bob at the inverse temperature $\beta$. Let Bob’s Hamiltonian be $H$  [^6] Since we do not care about the energy change of Alice’s system, it is not necessary to specify a Hamiltonian for Alice and we can also let the $|E_{x}\rangle_{A}$ to be any orthonormal set..
@@
We focus on teleportation protocols with rank-one projective measurements on $AA^{\prime}$, as general measurements with higher-rank Kraus operators have coarser resolution. Since we are interested in the trade-off relation as constrained by the entanglement resource, it is sufficient to consider measurements with resolution as sharp as possible  [^7] One could also ask how the resource of classical communication constrains the teleportations. Then it is also of interest to consider general measurements whose outcome consumes less communication resource to be sent.. For completeness, we discuss the case of general measurement instruments in Supplemental Materials. With rank-one measurement operators, the output states on $RB$ are pure so it is appropriate to use the entanglement entropy as a figure of merit for QIT. We prove the following result.
@@
We leave the proof in Supplemental Materials. This trade-off bound is tight. It is saturated by Bell measurements, and meanwhile Bob cannot extract energy. On the other hand, when Alice performs product measurements on $AA^{\prime}$ and measures the basis $\{|x\rangle_{A}\}$, Bob can extract the most amount of energy by applying the unitary that maps the steered state $|E_{x}\rangle_{B}$ to the ground state, and meanwhile quantum information cannot be teleported. The optimal QET, however, does not saturate the bound, leaving a gap of size $\log Z(\beta)$ (cf. Fig. 3 and discussions later).
There are some variants of this trade-off relation. Let $\rho_{B}:=\sum_{x}p_{x}\rho_{x}$ be the average output state of the teleportation protocol. Then we have
$E_{C}(\rho)+\beta E\leq E_{f}(\rho)+\beta E\leq S(B)_{\tau_{\beta}}\ ,$ (5)
@@
We leave the proof in Supplemental Materials. We reiterate that since an $E^{\mathrm{reg}}(\rho_{B})$ amount of energy can be extracted without any teleportation protocol in the asymptotic regime, it is only fair to subtract this contribution from the energy in (9). Only copies of the thermal state remain passive for all $n$, and they are called completely passive [47, 48]. For them, the remainder term (8) vanishes and (9) reduces to the previous result (4).
@@
Acknowledgments. We are grateful to Zhenbin Yang and Douglas Stanford for very helpful comments and discussions on traversable wormholes. We also thank Adam Brown, Raphael Bousso, Patrick Hayden, Henry Lin, Jingru Lu, and Jianming Zheng for the discussions.
Supplemental Materials
I Double trace Deformation as LOCC
@@ IV Proof of Theorem 2
V Teleportation schemes with general measurements.

### v2

@@ Abstract
Quantum energy teleportation (QET) is the phenomenon in which locally inaccessible energy is activated as extractable work through collaborative local operations and classical communication (LOCC) with an entangled partner. It closely resembles the more well-known quantum information teleportation (QIT) where quantum information can be sent through an entangled pair with LOCC. It is tempting to ask how QET is related to QIT. Here we report a first study of this connection. Despite the apparent similarity, we show that these two phenomena are not only distinct but moreover are mutually competitive. We show a perturbative trade-off relation between their performance in a thermal entangled chaotic many-body system, in which both QET and QIT are simultaneously implemented through a traversable wormhole in an emergent spacetime. Motivated by this example, we study a generic setup of two entangled qudits and prove a universal non-perturbative trade-off bound. It shows that for any teleportation protocol, the overall performance of QET and QIT together is constrained by the entanglement resource. We discuss some explanations of our results.
@@
To answer this question, we examine a setup where QIT and QET are simultaneously implemented with a single protocol. It concerns two entangled strongly coupled many-body quantum systems whose complicated dynamics can be more easily understood with the help of an emergent spacetime [10, 11, 12]. For a thermal field double state entangling two such systems, a special family of information teleportation protocols can be described through the physics of a traversable wormhole in the emergent spacetime [13, 14]. This has inspired a series of works exploring the deep connection between teleportation and wormholes [15, 16, 17, 18, 19, 20, 21, 22, 23]. We observe that the same protocol also simultaneously teleports energy. Interestingly, we find QET and QIT are competitive with each other in this example. Their exclusivity is manifest in the perturbative trade-off relation that we demonstrate.
Motivated by this observation, we look for universal non-perturbative trade-off relations between QET and QIT. To this end, we resort to the standard finite-dimensional setup that consists of two arbitrarily entangled systems with some arbitrary non-interacting Hamiltonian. We prove a universal non-perturbative trade-off bound for both the thermal-entangled states (Theorem 1) and arbitrary entangled states (Theorem 2), showing that the overall performance of QET and QIT altogether is upper-bounded by the amount of shared entanglement. The bounds are universal and non-perturbative in the sense that they apply to any pure entangled resource states, any measurement schemes of Alice and any conditional unitaries of Bob. The two theorems consitute the main results of this paper.
A motivating example: Traversable wormholes.—We find that a setup that naturally incorporates both teleportations is the traversable wormhole in holographic systems [13, 14]. It was discovered to be a convenient gravitational description of QIT. We use a concrete example of such systems in AdS2/CFT1 correspondence [14] to explain the result. The bulk theory is described by Jackiw–Teitelboim (JT) gravity. The boundary theory is described Schwarzian quantum mechanics [^1] Technically, Schwarzian quantum mechanics has a continuous energy spectrum. However, the Schwarzian describes the low energy dynamics of the Sachdev-Ye-Kitaev (SYK) model at the large $N$ limit [10], so we are allowed to think of the boundary dual of JT gravity as a normal quantum mechanical system like the SYK model with $N$ fermions., such as the Sachdev-Ye-Kitaev (SYK) model [10].
We consider a two-sided black hole with inverse temperature $\beta$, dual to the thermal field double state [24],
@@
The wormhole in Fig. 1 is a priori not traversable, as shown by the diagonal dotted causal wedges in the Penrose diagram. Any two points between the left and right causal wedges are space-like separated. On the boundary, this indicates an entangled state alone can not be used for communication. However, seminal works [13, 14] point out that the wormhole can be made traversable by introducing a double trace deformation on two sides $e^{igV_{AB}}$, where $V_{AB}:=\frac{1}{K}\sum_{i=1}^{K}O_{i}^{A}O_{i}^{B}$. This interaction couples $K$ pairs of simple operators $O_{i}$ on $\mathcal{O}(1)$ of fermions and $g$ controls the strength. The deformation with $g>0$ creates an ingoing shock wave with negative null energy that moves the black hole horizons inward, rendering a traversable wormhole [^2] For this description to be accurate, one technically chooses $K$ to be large $K\sim\mathcal{O}(N)$ so as to have a sharp bulk dual as a classical shock-wave geometry.. Namely, a particle carrying information thrown in the past from Alice’s side can now make its way to Bob’s side.
As far as Bob’s system is concerned, one can view the effect of the double trace deformation $e^{igV_{AB}}$ on Bob’s system, as an LOCC protocol $\prod_{x}\Pi_{x}^{A}\otimes U^{B}_{x}$, in which Alice measures her share of the entanglement pair and Bob implements a conditional unitary on his share upon receiving the measurement result from Alice after a time delay of $t$ (cf. Fig. 1 and Appendix A for details). Usually, $e^{igV_{AB}}$ is acted on the time-symmetric slice so $t=0$, but we consider acting it asymmetrically to fit the operational interpretation as an LOCC, so a delay of $t>0$ is inevitable. In this protocol, Alice can transmit information by inserting a simple operator with a species index at around the scrambling time $T$ before the measurement. The information will be transmitted to Bob’s side after Bob’s unitary decoding. After a scrambling time, the information will re-focus onto simple operators. (cf. [17] for details in the context of the SYK model.)
The first observation we make is that the traversable wormhole protocol actually also implements quantum energy teleportation. Without Alice’s communication, Bob cannot extract energy by doing a unitary, because Bob’s reduced density matrix is a thermal density matrix. With Alice’s measurement result $x$, Bob applies the conditional unitary decoder $U_{x}$ after some communication delay $t$. By energy conservation, Bob then gains energy that amounts to the negative energy that is injected into his system via the conditional unitary. In Appendix B, we calculate this teleported energy,
@@
We can make this trade-off more precise by quantifying the teleported information. We use $k$ different particle species indexed by $x\in[k]$ to carry the message. To calculate how much information can be teleported, we consider sending a maximally mixed message. By linearity, the state that Bob receives is $\rho=\frac{1}{k}\sum_{x=1}^{k}\rho_{x}$, where $\rho_{x}$ is the protocol output for the message $x$. The entropy of the message is measured by $S:=S(B)_{\rho}-S(B)_{\rho_{0}}$, where $\rho_{0}$ is Bob’s density matrix after the double trace deformation but without the message particle insertion [^3] One could use other measures such as the coherent (Holevo) information to quantify the quantum (classical) information capacity. They all behave very similarly to the subtracted entropy w.r.t. varying k [26, 27]. We therefore use the simplest figure of merit as a proof of principle..
We can calculate each term in $S$ using the (quantum) Ryu-Takayanagi formula [28, 29]. The state $\rho_{0}$ describes a bulk without matter and its entropy is given by $S(B)_{\rho_{0}}=A/4G_{N}(=\beta\langle H_{B}\rangle_{\rho_{0}})$, where $A$ is the area of the shifted bifurcation surface after the deformation, and it follows from the thermality of $\rho_{0}$ [^4] One can show this using the generalized gravitational entropy [30, 29]. On the other hand, $S(B)_{\rho}=A/4G_{N}+S_{\mathrm{matter}}$ [^5] Here the area term will be the same as for $\rho_{0}$ under the approximation of the messenger particle being light., where $S_{\mathrm{matter}}$ is the quantum field theoretic entropy of the message particle. We know that $S_{\mathrm{matter}}$ of a uniform mixture of $k$ particle species grows like $\log k$ until it saturates at the Bekenstein bound set by $\beta E$, and $E:=\langle H_{B}\rangle_{\rho}-\langle H_{B}\rangle_{\rho_{0}}$. Hence, $S$ increases with $\log k$, but eventually caps off due to the Bekenstein bound [25, 31, 32, 33]. The maximal $S$ thus measures the capacity of QIT [^6] This calculation is explicitly done using the replica trick in the upcoming paper by Lu-Yang-Zheng [34]..
Therefore, the loss in QET is totally compensated by the gain in QIT [^7] As always in thermodynamics, the (inverse) temperature $\beta$ here is the conversion factor that is necessary for comparing the distinct resources of information and energy.. In other words, upon varying the relevant parameters of the protocol, such as the conformal weight the message particle, the total variation of the combined QET+QIT figure of merit $S+\beta E$ is zero,
$\delta(S+\beta E)=\delta S+\beta\delta E=0\ .$ (3)
@@
There are some drawbacks of the trade-off (3). It is only perturbative because we cannot vary $m$ too much, as it causes significant backreaction rendering our semiclassical analysis inaccurate. Moreover, it only applies to a specific family of LOCC schemes that admit the bulk description as traversable wormholes. In seeking of a non-perturbative trade-off relation that applies to all LOCC protocols, we resort to a generic finite-dimensional setup.
Main results: Trade-off between QET and QIT.—We will now prove two trade-off bounds for the thermal field double state and arbitrary pure entangled states respectively. We first consider a finite-dimensional version of (1), $|\tau_{\beta}\rangle_{AB}$, between Alice and Bob at the inverse temperature $\beta$. Let Bob’s Hamiltonian be $H$ [^9] Since we do not care about the energy change of Alice’s system, it is not necessary to specify a Hamiltonian for Alice and we can also let the $|E_{x}\rangle_{A}$ to be any basis..
@@
We focus on teleportation protocols with rank-one projective measurements on $AA^{\prime}$, as general measurements with higher-rank Kraus operators have coarser resolution. Since we are interested in the trade-off relation as constrained by the entanglement resource, it is sufficient to consider measurements with resolution as sharp as possible [^10] One could also ask how the resource of classical communication constrains the teleportations. Then it is also of interest to consider general measurements whose outcome consumes less communication resource to be sent.. For completeness, we discuss the case of general measurement instruments in Appendix E. With rank-one measurement operators, the output states on $RB$ are pure so it is appropriate to use the entanglement entropy as a figure of merit for QIT. We prove the following result [^11] There are some variants of the trade-off relation (4) that we discuss in Appendix F..
@@
We leave the proof in Appendix C. This trade-off bound is tight. It is saturated by Bell measurements, and meanwhile Bob cannot extract energy. On the other hand, when Alice performs product measurements on $AA^{\prime}$ and measures the basis $\{|x\rangle_{A}\}$, Bob can extract the most amount of energy by applying the unitary that maps the steered state $|E_{x}\rangle_{B}$ to the ground state, and meanwhile quantum information cannot be teleported. The optimal QET, however, does not saturate the bound, leaving a gap of size $\log Z(\beta)$ (cf. Fig. 3 and discussions later).
@@
We leave the proof in Appendix D. We reiterate that since an $E^{\mathrm{reg}}(\rho_{B})$ amount of energy can be extracted without any teleportation protocol in the asymptotic regime, it is only fair to subtract this contribution from the energy in (8). Only copies of the thermal state remain passive for all $n$, and they are called completely passive [44, 45]. For them, the remainder term (7) vanishes and (8) reduces to the previous result (4).
@@
Acknowledgments.—We are grateful to Zhenbin Yang and Douglas Stanford for very helpful comments and discussions on traversable wormholes. We also thank Adam Brown, Raphael Bousso, Patrick Hayden, Henry Lin, Jingru Lu, and Jianming Zheng for the discussions. JW would like to thank the support from AFOSR (award FA9550-19-1-0369).
Appendix A Double trace Deformation as LOCC
@@ Appendix D Proof of Theorem 2
Appendix E Teleportation schemes with general measurements.
@@
Appendix F Variants of the trade-off relation (4)
There are some variants of our trade-off relation (4). Let $\rho_{B}:=\sum_{x}p_{x}\rho_{x}$ be the average output state of the teleportation protocol. Then we have
$E_{C}(\rho)+\beta E\leq E_{f}(\rho)+\beta E\leq S(B)_{\tau_{\beta}}\ ,$ (51)